A soda can contains about 14 grams of aluminum (Al), the most abundant element in Earth's crust. How many soda cans can be made from 52 moles of Al?


1
Expert's answer
2022-02-16T04:54:04-0500

"Moles = mass\/ rfm"

Rfm= 27

Mass= 27× 52= 1404g

Cans= 1404/14= 100